For all parts of this problem, let S be a finite set, with |S|=n.
(a) If f:S -> S is onto, does it necessarily follow that f is invertible? Why or why not?
(b) Compute the number of distinct functions g:S -> S that are onto.
(c) Find the number of functions f:S -> S that are not invertible.
1
Expert's answer
2020-12-08T06:43:48-0500

Let "S" be a finite set with "|S|=n".


(a) If "f:S \to S" is onto and "S" is finite, then "f" is injection. Indeed, suppose that "f" is not injection, then "f(a)=f(b)" for some "a\ne b". It follows that "|f(S)|\le n-1", and therefore "f(S)\subsetneq S". This contadicts to the fact that "f" is onto. Consequently, "f:S \to S" is bijection, and thus it is invertible.


(b) The number of distinct functions "g:S \to S" that are onto is equal to the number of all distinct bijection on the set "S". Taking into account that the number of all bijection on the set "S" is equal to the number of all permutation of the "n"-element set "S", we conclude that the number of distinct functions "g:S \to S" that are onto is equal to "n!."


(c) Let us compute the number of all functions "f:S \to S". For each "a\in S" the value "f(a)" of the function "f" at a point "a" can independently acquire one of the "n=|S|" values of the set "S". So, by the combinatorial rule of product, the number of all functions "f:S \to S" is equal to "\underbrace{n\cdot...\cdot n}_n=n^n".

Consequently, the number of functions "f:S \to S" that are not invertible is equal to "n^n-n!." .